# Intermittent thermal convection in jammed emulsions

### v1

@@ IV Conclusions
We studied the long-time behavior of a typical YSM, i.e., a jammed emulsion, under thermal convection. We pointed out a novel phenomenon, never reported before, whereby convection occurs through fluidization-rigidity transitions, entailing heat flux bursts occurring in an otherwise conductive background. Once thermal convection is initiated, temperature field mixing causes stress fluctuations leading to a halting of convection if the stress drops below the yield value. Leveraging on numerical simulations [46], we can observe that, remarkably, residual plastic activity persists even during macroscopic resting periods, thus acting as a latent trigger for subsequent convective events. This plasticity stems from the inherently disordered structure of the emulsion, resulting in an intermittent onset of convective transport without discernible regularity, which cannot be predicted based on continuum equations coupled with bulk rheology. Fluidization events and heat bursts are associated with droplets’ coalescence, eventually leading to a phase-inverted emulsion exhibiting spatial heterogeneity: to our knowledge, such a final state remains unexplored and warrants further investigation, both morphologically and rheologically. For a given emulsion, we studied the intermittency phenomenon at a fixed Rayleigh number ($\operatorname{Ra}$), and we expect that it can only survive in a specific range of $\operatorname{Ra}$. Indeed, for very low $\operatorname{Ra}$ values, convection is suppressed, whereas for large $\operatorname{Ra}$ values a quicker transition to the convective state in the phase-inverted emulsion is prompted through more frequent coalescence events, bypassing the intermittent behavior.
@@ Acknowledgements
This work has been carried out within the TEXTAROSSA project (G.A. H2020-JTI-EuroHPC-2019-1 No. 956831). MS and MB acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 - CUP E83C22003230001, Mission 4 Component 2 Investment 1.4, funded by the European Union - NextGenerationEU.

### v2

@@ IV Conclusions
We studied the long-time behavior of a typical YSM, i.e., a jammed emulsion, under thermal convection. We pointed out a novel phenomenon, never reported before, whereby convection occurs through fluidization-rigidity transitions, entailing heat flux bursts occurring in an otherwise conductive background. Once thermal convection is initiated, temperature field mixing causes stress fluctuations leading to a halting of convection if the stress drops below the yield value. Leveraging on numerical simulations [48], we can observe that, remarkably, residual plastic activity persists even during macroscopic resting periods, thus acting as a latent trigger for subsequent convective events. This plasticity stems from the inherently disordered structure of the emulsion, resulting in an intermittent onset of convective transport without discernible regularity, which cannot be predicted based on continuum equations coupled with bulk rheology. Fluidization events and heat bursts are associated with droplets’ coalescence, eventually leading to a phase-inverted emulsion exhibiting spatial heterogeneity: to our knowledge, such a final state remains unexplored and warrants further investigation, both rheologically and morphologically. For a given emulsion, the intermittency phenomenon could survive in a finite range of values of the Rayleigh number $\operatorname{Ra}$. Indeed, for values of $\operatorname{Ra}$ below this range, convection is suppressed, whereas, for larger values, phase inversion is observed without transient intermittency.
@@ Acknowledgements
This work has been carried out within the TEXTAROSSA project (G.A. H2020-JTI-EuroHPC-2019-1 No. 956831). FP, MS and MB acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP E83C22003230001 and CUP B93C22000620006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.